OpenStudy (anonymous):

the price of a radio changes from $40.00 to $44.00. A. 10% decrease B.10% increase C.36.4% decrease D. 40% increase

OpenStudy (anonymous):

Can you tell me if it was an increase or a decrease?

OpenStudy (anonymous):

increase

OpenStudy (anonymous):

alright so you know its either b or d

OpenStudy (anonymous):

so do you know how to calculate the percent increase?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well.....we are measuring the change in something right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so when we are measuring the change in something that means we usually will subtract two things. What do you think we will subtract?

OpenStudy (anonymous):

44.00-40.00 right ??

OpenStudy (anonymous):

yes so the new price minus the initial price.

OpenStudy (anonymous):

Now its also called the percentage change in something

OpenStudy (anonymous):

to change something to a percent we divide by something and multiply by 100

OpenStudy (anonymous):

what do you think we should divide by?

OpenStudy (anonymous):

i got 4 when i subtracted

OpenStudy (anonymous):

4

OpenStudy (anonymous):

you should have gotten 4 when you subtracted but now we have to convert that to a percent

OpenStudy (anonymous):

so what do you think we should divide by?

OpenStudy (anonymous):

@katalina13

OpenStudy (anonymous):

100

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

we divide by the initial price

OpenStudy (anonymous):

40

OpenStudy (anonymous):

im sorry im not good at this

OpenStudy (anonymous):

right then we multiply by 100

OpenStudy (anonymous):

don't worry I'm here to help you not judge you :)

OpenStudy (anonymous):

i got 1,100

OpenStudy (anonymous):

not quite remember its this \[\frac{ (44-40 }{ 40 }*100\]

OpenStudy (anonymous):

opps forgot my close bracket on top :P

OpenStudy (anonymous):

i got what i did wrong now i divide by 44

OpenStudy (anonymous):

1,000

OpenStudy (anonymous):

ummm definitely not 1000

OpenStudy (anonymous):

what is \[\frac{ 4 }{ 40 }\] ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

not quite its \[\frac{ 1 }{ 10 }\]

OpenStudy (anonymous):

or 0.1

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

if it were \[\frac{ 40 }{ 4 }\] then it would be 10

OpenStudy (anonymous):

so now that we know its 0.1 what is 0.1*100?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

or is my calculator messed up

OpenStudy (anonymous):

so what is the answer?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

thanks so much for your help

OpenStudy (anonymous):

not a problem don't forget to close the question.

OpenStudy (anonymous):

ok